OpenStudy (anonymous):

what is the simplified sum of √3+√12?

OpenStudy (anonymous):

3 rad3

OpenStudy (anonymous):

\[\sqrt{3}+2\sqrt{3}=3\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{3}+\sqrt{12}=\sqrt{3}+2\sqrt{3}=3\sqrt{3}\]